OpenStudy (anonymous):

An object is in simple harmonic motion. Find an equation for the motion given that the period is pi/4 and, at time t=o, y=1, y'=0. What is the amplitude? What is the frequency?

OpenStudy (anonymous):

you will need the formulas for this

OpenStudy (anonymous):

Period is basically the reciprocal of freqiency

OpenStudy (anonymous):

yes, i believe we will use \[y=Asin(\omega t+\phi _{0})\]

OpenStudy (anonymous):

and amplitude = -angular frequency^2 times x

OpenStudy (anonymous):

\[\omega = 8\]

OpenStudy (anonymous):

I'll list the formulas you will need. remember them

OpenStudy (anonymous):

\[\omega = 2 \pi f\] \[a = -\omega ^{2} x\] \[f = \frac{ 1 }{ T }\]

OpenStudy (anonymous):

yes and I have f = 4pi and i have omega as 8.. but how to i find the rest.

OpenStudy (anonymous):

thr amplitude has to be infered

OpenStudy (anonymous):

y'=0 means a minima or a maxima (i.e the amplitude)

OpenStudy (anonymous):

therefore, the question stated the amplitude

OpenStudy (anonymous):

which is 1

OpenStudy (anonymous):

so you have the amplitude and frequency. Find the angular frequency and use the SHM formula to get the equation

OpenStudy (anonymous):

so you have the amplitude and frequency. Find the angular frequency and use the SHM formula to get the equation\[x = x _{o} \sin (\omega t)\]

OpenStudy (anonymous):

okay so then i have \[y=\sin(8t+\phi _{0})\] how do i find phi

OpenStudy (anonymous):

\( t=0, y=1, y'=0\), it says that \(t=0\) is a critical point meaning \(y\) is at its maximum so amplitude is \(1\).

OpenStudy (anonymous):

Now \(\sin(0) = 0\) and \(\cos(0)=1\) which tells you that you want to use \(\cos\) rather than \(\sin\), in which case your initial angle is just \(0\).

OpenStudy (anonymous):

the answer key says phi is pi/2, i do not know how they got that... so the equation is \[y=\sin(8t+\frac{ \Pi }{ 2 })\]

OpenStudy (anonymous):

They likely used co-function identities. \[ \cos(t) = \sin(\pi/2-t) \]

OpenStudy (anonymous):

It is also known that \(\sin(\pi/2\))=1

OpenStudy (anonymous):

well the class is differential equation, so i am certain that i have to use the derivative of the general equation somehow.

OpenStudy (anonymous):

Meaning if you start your initial angle at \(\pi/2\) then you'll start at your amplitude.